OpenStudy (anonymous):

how do I write this with positive exponents? z^-2 ----- z^-3

OpenStudy (anonymous):

would it be 1 ____ Z

OpenStudy (anonymous):

\[\Large x^{-\color{red}a}=\frac{1}{x^\color{red}a} \]

OpenStudy (anonymous):

ok so would it be \[1\]

OpenStudy (anonymous):

1 over Z

OpenStudy (anonymous):

\[\Large \frac{z^{-2}}{z^{-3}} = \frac{1}{z^{2}z^{-3}} = \frac{1}{z^{-1}} = \frac{1}{\frac{1}{z}}=z \]

OpenStudy (anonymous):

ok great thank you!